# Hierarchical Diffusion Motion Planning with Task-Conditioned Uncertainty-Aware Priors

### v1

@@ Abstract
We propose a novel hierarchical diffusion planner that embeds task and motion structure directly in the noise model. Unlike standard diffusion-based planners that use zero-mean, isotropic Gaussian noise, we employ a family of task-conditioned structured Gaussians whose means and covariances are derived from Gaussian Process Motion Planning (GPMP): sparse, task-centric key states or their associated timings (or both) are treated as noisy observations to produce a prior instance. We first generalize the standard diffusion process to biased, non-isotropic corruption with closed-form forward and posterior expressions. Building on this, our hierarchy separates prior instantiation from trajectory denoising: the upper level instantiates a task-conditioned structured Gaussian (mean and covariance), and the lower level denoises the full trajectory under that fixed prior. Experiments on Maze2D goal-reaching and KUKA block stacking show improved success rates, smoother trajectories, and stronger task alignment compared to isotropic baselines. Ablation studies indicate that explicitly structuring the corruption process offers benefits beyond simply conditioning the neural network. Overall, our method concentrates probability mass of prior near feasible, smooth, and semantically meaningful trajectories while maintaining tractability. Our project page is available at https://hta-diffusion.github.io.
@@ I Introduction
In this work, we propose to encode such structure in the form of a task-aware noise model. This design enables our diffusion planner to respect smoothness and constraint satisfaction by construction. We realize this idea with a two-level hierarchical diffusion framework. The upper level uses a denoising diffusion probabilistic model (DDPM) [9] with isotropic Gaussian noise to produce sparse, task-centric key states. For example, in navigation tasks, key states can be chosen as evenly spaced waypoints between the start and goal. In manipulation tasks, they naturally align with task phases (e.g., approach, grasp, lift), augmented by a binary contact-flag trajectory. [^1] Hereafter, we use key states to denote both waypoints in navigation and keyframe states in manipulation. We treat upper-level key states as soft observations with explicit uncertainty, so these imperfect predictions guide sampling without hard constraints or rejection. The lower-level diffusion then generates the full trajectory under a task-conditioned Gaussian, obtained by conditioning a linear time-varying Gaussian process (LTV-GP) motion prior on the key states, inspired by Gaussian process motion planning (GPMP) [10]. This yields a non-isotropic and temporally correlated noise model whose mean and covariance encourage smooth dynamics between key states, while the key states act as anchors that prevent global oversmoothing and allow the trajectory to adapt at sharp turns or task-phase changes. Accordingly, this framework produces temporally consistent and constraint-aware plans.
@@
Prior diffusion works typically keep the corruption isotropic and inject structural constraints through conditioning or guidance during inference [8, 1, 2]. For instance, conditioning has been used to incorporate visual context from observations or to inject goal information [3, 6]. Guidance mechanisms and auxiliary objectives incorporate dynamics, task rewards, or safety specifications into the generative process [11, 12]. In addition to conditioning and guidance, task-centric representations have been leveraged to encode constraints, for example object-centric demonstrations [13] or SE(3)-based cost fields that couple grasp and motion optimization [14]. Projection-based methods have also been explored, where diffusion samples are mapped into feasible sets such as dynamically admissible trajectories [15] or collision-free multi-robot paths [16]. Beyond these categories, diffusion models have also been integrated with model-based motion planning [17] and with neural dynamics models [18].
@@
Structured corruptions during training have been explored primarily outside robotics and remain rare within it. Examples include non-Gaussian but zero-mean choices [19], per-pixel variance schedules without temporal or robotics priors [20], and non-isotropic formulations that capture joint dependencies in human pose but still assume zero-mean noise and omit temporal correlations along the horizon, limiting their use for trajectory state conditioning [21]. Mixture priors have also been proposed for multimodality and controllability, where Gaussian mixture parameters may be predefined through clustering [22] or heuristic structure [23], learned end-to-end to adaptively cover target support [24], or dynamically adjusted during inference by solver-based approaches [25]. In robotics, structured corruptions remain rare, with only a few recent efforts—for example, integrating coarse trajectory predictions so that diffusion acts as a refiner [26], introducing temporally correlated noise such as colored Gaussian perturbations [27], or employing diffusion bridges that initiate denoising from informative prior actions in navigation tasks [28]. While these studies highlight the benefits of structured corruptions for distributional expressiveness, they do not leverage structured priors to encode task conditions and temporal smoothness simultaneously. In contrast, our design directly biases the generative process toward trajectories that are both dynamically consistent and task-aware, rather than only improving multi-modality.
@@
Hierarchical diffusion has been used to separate coarse-level decisions from finer-level motion, for example: subgoal generation followed by refinement [29], cascaded global and local prediction with online repair [30], progressive coarse-to-fine refinement processes that achieve real-time planning frequencies [31], contact-guided manipulation policies that decompose high-level contact planning from low-level trajectory generation [32], kinematics-aware frameworks with high-level end-effector prediction and low-level joint generation [33], and diffusion–flow hybrids that pair high-level diffusion with low-level rectified flow [34]. A common pattern feeds high-level outputs to the low level as conditioning, while low-level corruption remains isotropic. This leaves the forward process and loss unchanged, so temporal coupling, boundary consistency, and waypoint fidelity must be relearned at every step. Our hierarchy instead reparameterizes the low-level prior by conditioning a GP on upper-level key states. This yields a trajectory posterior whose mean interpolates the key states and whose covariance couples time steps, so reverse updates follow a Mahalanobis geometry that narrows the search to trajectories consistent between key states and treats imperfect key states probabilistically. This improves feasibility without post hoc constraints or rejection.
@@
We evaluate on Maze2D goal reaching and KUKA block stacking from the Diffuser benchmarks [2]. We compare (i) a single layer diffuser with isotropic noise, (ii) a hierarchical diffuser that only conditions on upper-level outputs while keeping isotropic noise, and (iii) our full hierarchical model that constructs the structured prior from upper-level key states and denoises under that prior. Across tasks, our model (iii) outperforms the conditioning-only baseline (ii), indicating that conditioning alone is insufficient. Embedding structure in the generative noise yields consistent gains in success and collision or constraint satisfaction.
@@
  • A two-level hierarchical diffusion planner where the upper-level DDPM infers sparse, task-centric key states and the lower level denoises under a GPMP-based motion prior conditioned on these states. Key states are treated as soft observations with explicit uncertainty, encoding smoothness, boundary and phase consistency, and handling non-stationarities without hard constraints.
  • A diffusion formulation with task-conditioned non-isotropic Gaussian corruption obtained by conditioning a GPMP-based prior on key states. We derive closed-form forward marginals and posteriors for variational training, yielding a Mahalanobis loss that generalizes MSE while preserving the standard DDPM pipeline.
  • A conceptual and empirical comparison to conditioning-only diffusion hierarchies. Reparameterizing the low-level prior alters the diffusion landscape and concentrates sampling on trajectories consistent with key states, improving feasibility without extra constraints or loss terms.
  • Experiments on Maze2D goal reaching and KUKA block stacking showing higher success and better constraint satisfaction than baselines with isotropic noise or conditioning alone, with ablations demonstrating the benefits of the structured prior.
Figure 1: Overview of our hierarchical diffusion framework. Training (left): From dataset trajectories we extract task-relevant key states and timings $(Y,C)$. These supervise the upper-level diffusion (Alg. 1) and also define the structured prior for the lower level, trained under this task-aware noise model (Alg. 2). Test-time planning (right): For a new task specification, the upper level provides $(Y,C)$ via rule-based heuristics or the learned diffusion. These outputs instantiate the task-specific prior, from which the lower level initializes and denoises to generate a full trajectory (Alg. 3).
@@ II-A DDPM for motion planning
Let a robot trajectory be $\tau=[x_{1},\ldots,x_{H}]\in\mathbb{R}^{H\times d}$, where $x_{t}\in\mathbb{R}^{d}$ is the robot state at time step $t$ and $H$ is the planning horizon. We denote by $\tau^{i}$ the trajectory after $i$ forward diffusion steps.
@@
The one-step forward posterior is also Gaussian,
@@ II-B Gaussian Process Motion Planning
$\tau(t)\sim\mathcal{GP}(\tilde{\mu}(t),\tilde{\mathcal{K}}(t,t^{\prime})),$ (4)
whose mean and kernel are induced by a linear time varying (LTV) dynamics prior [10]. Consider the stochastic LTV model $\dot{x}(t)=A(t)x(t)+u(t)+w(t)$ where $w(t)$ is zero mean white noise with spectral density $Q_{C}\succeq 0$. Let the initial mean and covariance at time $t_{0}$ be $\mu_{0}$ and $\mathcal{K}_{0}$, and let $\Phi(t,s)$ denote the state transition matrix of $A(\cdot)$. Then the GP has mean $\tilde{\mu}(t)=\Phi(t,t_{0})\mu_{0}+\int_{t_{0}}^{t}\Phi(t,s)u(s)\,ds$ and kernel $\tilde{\mathcal{K}}(t,t^{\prime})=\Phi(t,t_{0})\mathcal{K}_{0}\Phi(t^{\prime},t_{0})^{\top}+\int_{t_{0}}^{\min\{t,t^{\prime}\}}\Phi(t,s)Q_{C}\Phi(t^{\prime},s)^{\top}ds$, which admit sparse and exactly computable discrete counterparts.
@@
Figure 2: Trajectory generation pipeline. Input: task specification (e.g., Maze2D goal reaching; KUKA grasp-and-place) and initial conditions. Upper level produces key states (stack of which is denoted with $Y$) and their timings (equivalent to selection matrix $C$), either by diffusion module and/or rules. For Maze2D, they are waypoints with fixed timings. For KUKA, they are grasp/release states and contact flags, which implicitly imply timing. With $(Y,C)$, we construct a Gaussian prior by conditioning a GP over trajectories, yielding $(\mu,\mathcal{K})$ via (9)–(10). The shaded regions illustrate covariance structure $\mathcal{K}$ around the trajectory, showing tighter variance near key states and looser variance elsewhere. The lower level then generates the full trajectory by reverse diffusion from $\mathcal{N}(\mu,\mathcal{K})$.
@@ III Method
III-A Diffusion with a task conditioned Gaussian and hierarchical planner overview
Let $\mathcal{D}$ be a dataset of robot trajectories where each clean sample $\tau\!\in\!\mathcal{D}$ is paired with task information that may admit multiple feasible solutions. In our settings, this task information includes start and goal or a waypoint template in Maze2D, and grasp and place timing with key joint configurations in KUKA manipulation block stacking.
It is desirable for the corruption (forward) process to retain task information so that reverse diffusion starts from a task-relevant noise prior. To this end, we modify the standard DDPM to use a non-isotropic Gaussian with a task conditioned mean and covariance. Reusing the schedule notation from Sec. II-A, with $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{i}=\prod_{k=1}^{i}\alpha_{k}$, we define
$\tau^{i}\sim\mathcal{N}\!\Big(\sqrt{\alpha_{i}}\,\tau^{i-1}+(1-\sqrt{\alpha_{i}})\,\mu,\;(1-\alpha_{i})\,\mathcal{K}\Big),$ (5)
where $\mu\!\in\!\mathbb{R}^{H\times d}$ and $\mathcal{K}\!\in\!\mathbb{R}^{Hd\times Hd}$ with $\mathcal{K}\succeq 0$ encode task-dependent trajectory structure. $\tau^{0}$ is just $\tau$.
@@ III-A Diffusion with a task conditioned Gaussian and hierarchical planner overview
$\tau^{i}\mid\tau^{0}\sim\mathcal{N}\!\Big(\sqrt{\bar{\alpha}_{i}}\,\tau^{0}+(1-\sqrt{\bar{\alpha}_{i}})\,\mu,\;(1-\bar{\alpha}_{i})\,\mathcal{K}\Big).$ (6)
The one step forward posterior is also Gaussian,
$q(\tau^{i-1}\!\mid\!\tau^{i},\tau^{0})=\mathcal{N}\!\big(\tau^{i-1};\,\tilde{\mu}_{i}(\tau^{i},\tau^{0},\mu),\,\tilde{\beta}_{i}\,\mathcal{K}\big),$ (7)
with coefficients $\tilde{\mu}_{i}$ and $\tilde{\beta}_{i}$ that directly generalize the DDPM posterior mean $\hat{\mu}_{i}$ and variance coefficient $\hat{\beta}_{i}$ from Sec. II-A (explicit derivations are given in the Appendix). As $N\to\infty$, the terminal distribution approaches $\tau^{N}\!\sim\!\mathcal{N}(\mu,\mathcal{K})$, so at test time we initialize the reverse diffusion from this task-specific Gaussian.
At test time, we aim to generate a trajectory $\tau$ that is both dynamically feasible and aligned with task constraints. Consequently, the mean $\mu$ and covariance $\mathcal{K}$ of the non-isotropic Gaussian should encode both dynamics and task information. We achieve this with a two-level hierarchy. The upper level extracts task information that is available without access to the full trajectory and uses it to instantiate $(\mu,\mathcal{K})$. Concretely, it generates sparse task-centric key states with associated uncertainties that we treat as soft observations. The lower level integrates these key states with the GPMP prior (4) from Sec. II-B. Conditioning the GP prior on the soft observations yields a posterior mean and covariance that define $\mu$ and $\mathcal{K}$ in (5). During test time, reverse diffusion then proceeds under this non-isotropic and time correlated Gaussian, starting from $\mathcal{N}(\mu,\mathcal{K})$.
@@
The overall planner therefore has two functions: the upper level produces key states and their associated timings, and the lower level generates the full trajectory by reverse diffusion given $(\mu,\mathcal{K})$. Inputs are a task specification and an initial condition. An overview of the framework is shown in Fig. 1. The next subsections detail each level.
III-B Upper level: key states and timings

### v2

@@ Abstract
We propose a novel hierarchical diffusion planner that embeds task and motion structure directly into the noise model. Unlike standard diffusion-based planners that rely on zero-mean, isotropic Gaussian corruption, we introduce task-conditioned structured Gaussians whose means and covariances are derived from Gaussian Process Motion Planning (GPMP), explicitly encoding trajectory smoothness and task semantics in the prior. We first generalize the standard diffusion process to biased, non-isotropic corruption with closed-form forward and posterior expressions. Building on this formulation, our hierarchical design separates prior instantiation from trajectory denoising. At the upper level, the model predicts sparse, task-centric key states and their associated timings, which instantiate a structured Gaussian prior (mean and covariance). At the lower level, the full trajectory is denoised under this fixed prior, treating the upper-level outputs as noisy observations. Experiments on Maze2D goal-reaching and KUKA block stacking show consistently higher success rates and smoother trajectories than isotropic baselines, achieving dataset-level smoothness substantially earlier during training. Ablation studies further show that explicitly structuring the corruption process provides benefits beyond neural conditioning the denoising network alone. Overall, our approach concentrates the prior’s probability mass near feasible and semantically meaningful trajectories. Our project page is available at https://hta-diffusion.github.io.
@@ I Introduction
In this work, we propose to encode such structure in the form of a task-aware noise model. This design enables our diffusion planner to respect smoothness and constraint satisfaction by construction. We realize this idea with a two-level hierarchical diffusion framework. The upper-level planner produces sparse, task-centric key states (or key constraints). We treat these key states as soft observations with explicit uncertainty, allowing imperfect predictions to guide sampling without hard constraints or rejection. The lower-level diffusion then generates the full trajectory under a task-conditioned Gaussian, obtained by conditioning a linear time-varying Gaussian process (LTV-GP) motion prior on the key states, inspired by Gaussian process motion planning (GPMP) [18]. This yields a non-isotropic, temporally correlated noise model whose mean and covariance encourage smooth dynamics between key states, while the key states act as anchors that prevent global oversmoothing and allow the trajectory to adapt at sharp turns or task-phase changes. Accordingly, this framework produces temporally consistent and constraint-aware plans.
@@
Previous diffusion works typically keep the corruption isotropic and inject structural constraints through conditioning or guidance during inference [21, 4, 12]. For instance, conditioning has been used to incorporate visual context from observations or to inject goal information [6, 32]. Guidance mechanisms and auxiliary objectives incorporate dynamics, task rewards, or safety specifications into the generative process [24, 31]. In addition to conditioning and guidance, task-centric representations have been leveraged to encode constraints, for example object-centric demonstrations [11] or SE(3)-based cost fields that couple grasp and motion optimization [26]. Projection-based methods have also been explored, where diffusion samples are mapped into feasible sets such as dynamically admissible trajectories [3] or collision-free multi-robot paths [15]. Beyond these categories, diffusion models have also been integrated with model-based plan [33] or with neural dynamics models [30].
@@
Structured corruptions during training have been explored primarily outside robotics. Examples include non-Gaussian but zero-mean choices [19], per-pixel variance schedules without temporal or robotics priors [34], and non-isotropic formulations that capture joint dependencies in human pose but still assume zero-mean noise and omit temporal correlations along the horizon, limiting their use for trajectory state conditioning [7]. Mixture priors have also been proposed for multimodality and controllability, where Gaussian mixture parameters may be predefined through clustering [13] or heuristic structure [28], learned end-to-end to adaptively cover target support [2], or dynamically adjusted during inference by solver-based approaches [9]. In robotics, structured corruptions remain rare, with only a few recent efforts—for example, integrating coarse trajectory predictions so that diffusion acts as a refiner [35], introducing temporally correlated noise such as colored Gaussian perturbations [14], or employing diffusion bridges that initiate denoising from informative prior actions in navigation tasks [22]. While these studies highlight the benefits of structured corruptions for distributional expressiveness, they do not leverage structured priors to encode task conditions and temporal smoothness simultaneously. In contrast, our design directly biases the generative process toward trajectories that are both dynamically consistent and task-aware, rather than only improving multimodality.
@@
Hierarchical diffusion has been used to separate coarse-level decisions from finer-level motion, for example: subgoal generation followed by refinement [5], cascaded global and local prediction with online repair [25], progressive coarse-to-fine refinement processes that achieve real-time planning frequencies [8], contact-guided manipulation policies that decompose high-level contact planning from low-level trajectory generation [27], kinematics-aware frameworks with high-level end-effector prediction and low-level joint generation [16], and diffusion–flow hybrids that pair high-level diffusion with low-level rectified flow [20]. A common pattern feeds high-level outputs to the low level as conditioning, while low-level corruption remains isotropic. This leaves the forward process and loss unchanged, while low-level corruption remains isotropic. Our hierarchy instead reparameterizes the low-level prior by conditioning a GP on upper-level key states. This yields a trajectory posterior whose mean interpolates the key states and whose covariance couples time steps, so reverse diffusion operates in a Mahalanobis geometry that restricts sampling to trajectories consistent between key states while modeling key-state uncertainty. This improves feasibility without post hoc constraints or rejection.
@@
We evaluate on Maze2D goal reaching and KUKA block stacking from the Diffuser benchmarks [12]. We compare (i) a single-layer diffuser with isotropic noise, (ii) a hierarchical diffuser with conditioning only, (iii) a diffuser with cost guidance at inference, and (iv) our full hierarchical model with a structured, key-state–conditioned prior. Across tasks, our model outperforms all the baselines, indicating that embedding structure in the generative noise yields consistent gains in task success over just conditioning or guidance.
@@
  • A two-level hierarchical diffusion planner where the upper-level planner infers sparse, task-centric key states and the lower-level planner denoises under a GPMP-based motion prior conditioned on them. Key states are treated as soft observations with explicit uncertainty, encoding smoothness while encouraging satisfaction of task constraints without imposing hard constraints.
  • A diffusion formulation with task-conditioned non-isotropic Gaussian corruption, obtained by conditioning a GPMP-based prior on key states, with closed-form forward marginals and posteriors for variational training. This yields a Mahalanobis loss that generalizes MSE while preserving the standard DDPM pipeline.
  • Experiments on Maze2D goal reaching and KUKA block stacking show higher success and better constraint satisfaction than isotropic and conditioning-only baselines, validating that reparameterizing the low-level prior reshapes the diffusion landscape and concentrates samples around key-state–consistent trajectories.
@@ II-A DDPM for motion planning
Let a robot trajectory be a stacked sequence of states $\tau=[x_{1};\ldots;x_{H}]\in\mathbb{R}^{H\cdot d}$, where $x_{t}\in\mathbb{R}^{d}$ is the robot state at time step $t$ and $H$ is the planning horizon. We denote by $\tau^{i}$ the trajectory after $i$ forward diffusion steps.
@@
The one-step reverse posterior is also Gaussian,
@@
Figure 1: Trajectory generation pipeline. Input: task goal and initial conditions. Upper level produces matrices $Y$ and $C$ for sparse key constraint $C\cdot\tau=Y$, either by diffusion module and/or rules. (e.g., if the constraints are simple sparse waypoints, $Y$ is a stack of key states and $C$ is a selection matrix, which corresponds to designation of timings for the key states.) With $(Y,C)$, we construct a Gaussian prior by conditioning a GP over trajectories, yielding $(\xi,\mathcal{K})$ via (8)–(9). The shaded regions illustrate covariance structure $\mathcal{K}$ around the trajectory, showing tighter variance near key states and looser variance elsewhere. The lower level then generates the full trajectory by reverse diffusion from $\mathcal{N}(\xi,\mathcal{K})$.
@@ II-B Gaussian Process Motion Planning
$x(t)\sim\mathcal{GP}(\tilde{\xi}(t),\tilde{\mathcal{K}}(t,t^{\prime})),$ (4)
whose mean and kernel are induced by a linear time varying (LTV) dynamics prior [18]. Consider the stochastic LTV model $\dot{x}(t)=A(t)x(t)+u(t)+w(t)$ where $u(t)$ is control input and $w(t)$ is zero mean white noise with spectral density $Q_{C}\succeq 0$. Let the initial mean and covariance at time $t_{0}$ be $\xi_{0}$ and $\mathcal{K}_{0}$, and let $\Phi(t,s)$ denote the state transition matrix of $A(\cdot)$. Then the GP has mean $\tilde{\xi}(t)=\Phi(t,t_{0})\xi_{0}+\int_{t_{0}}^{t}\Phi(t,s)u(s)\,ds$ and kernel $\tilde{\mathcal{K}}(t,t^{\prime})=\Phi(t,t_{0})\mathcal{K}_{0}\Phi(t^{\prime},t_{0})^{\top}+\int_{t_{0}}^{\min\{t,t^{\prime}\}}\Phi(t,s)Q_{C}\Phi(t^{\prime},s)^{\top}ds$, which admit sparse and exactly computable discrete counterparts.
@@ III Method
Let $\mathcal{D}$ be a dataset of robot trajectories where each clean sample $\tau^{0}\!\in\!\mathcal{D}$ is paired with task information that may admit multiple feasible solutions. It is desirable for the corruption (forward) process to retain task information so that reverse diffusion starts from a task-relevant noise prior. To this end, we modify the standard DDPM to use a non-isotropic Gaussian with a task conditioned mean and covariance. Reusing the schedule notation from Sec. II-A, with $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{i}=\prod_{k=1}^{i}\alpha_{k}$, we define
$\tau^{i}\sim\mathcal{N}\!\Big(\sqrt{\alpha_{i}}\,\tau^{i-1}+(1-\sqrt{\alpha_{i}})\,\xi,\;(1-\alpha_{i})\,\mathcal{K}\Big),$ (5)
where $\xi\!\in\!\mathbb{R}^{H\cdot d}$ and $\mathcal{K}\!\in\!\mathbb{R}^{Hd\times Hd}$ with $\mathcal{K}\succeq 0$ encode task-dependent trajectory structure.
@@
$\tau^{i}\mid\tau^{0}\sim\mathcal{N}\!\Big(\sqrt{\bar{\alpha}_{i}}\,\tau^{0}+(1-\sqrt{\bar{\alpha}_{i}})\,\xi,\;(1-\bar{\alpha}_{i})\,\mathcal{K}\Big).$ (6)
The one step backward posterior is also Gaussian,
$q(\tau^{i-1}\!\mid\!\tau^{i},\tau^{0})=\mathcal{N}\!\big(\tau^{i-1};\,\tilde{\mu}_{i}(\tau^{i},\tau^{0},\xi),\,\tilde{\beta}_{i}\,\mathcal{K}\big),$ (7)
where
$\tilde{\mu}_{i}(\tau^{i},\tau^{0},\xi):=\frac{\sqrt{\bar{\alpha}_{i-1}}\,\beta_{i}}{1-\bar{\alpha}_{i}}\,\tau^{0}+\frac{\sqrt{\alpha_{i}}\,(1-\bar{\alpha}_{i-1})}{1-\bar{\alpha}_{i}}\,\tau^{i}+\eta_{i}\xi,$
$\eta_{i}$ $:=\frac{1}{1-\bar{\alpha}_{i}}\left[\beta_{i}\bigl(1-\sqrt{\bar{\alpha}_{i-1}}\bigr)-\sqrt{\alpha_{i}}\,(1-\bar{\alpha}_{i-1})\bigl(1-\sqrt{\alpha_{i}}\bigr)\right],$
$\tilde{\beta}_{i}$ $=\left(\frac{\alpha_{i}}{1-\alpha_{i}}+\frac{1}{1-\bar{\alpha}_{i-1}}\right)^{-1}$
are generalization of the DDPM posterior mean and variance coefficient (derivations are given in the Appendix). As $N\to\infty$, the terminal distribution approaches
$\tau^{N}\!\sim\!\mathcal{N}(\xi,\mathcal{K}).$
We aim to generate a trajectory $\tau^{0}$ that is both smooth and aligned with task constraints. To this end, the mean $\xi$ and covariance $\mathcal{K}$ of the non-isotropic Gaussian are constructed using a two-level hierarchy based on GPMP (Sec. II-B). The GP kernel naturally encodes smoothness, while task-specific conditioning is incorporated through GP conditioning, as detailed in Sec. III-B.
@@
The overall planner consists of two components: the upper level produces task-relevant key states and their timings, which are used to construct the task-conditioned prior $(\xi,\mathcal{K})$; the lower level then generates the full trajectory by reverse diffusion under this fixed prior. The inputs are a task specification and an initial condition. The next subsections detail each level.
III-A Upper level: key states and timings